5. Calculate the pH of a buffer solution prepared by mixing 75 mL of 1.0 M lactic acid and 25 mL of a 1.0M sodium lactate. (pka of lactic acid 3.86) 6. Calculate the change in pH that occurs when 0.01moles of OH are added to a 1L buffered solution that contains 0.5M acetic acid (HC2H302) and 0.5M acetate ion (C2H302). Acetic acid pka = 4.76
Answer #1

5. Its an example for Acidic Buffer

pH = pKa + log (concentration x volume)salt/ (concentration x volume)Acid

= 3.86 + log [(1x 25)/ 1x 75)]

= 3.86 + log (1/3)

= 3.86 -0.48

pH = 3.38

6. Its an example for Acidic Buffer

The added OH- will reacts with acid and forms salt . Hence salt concentration will increases and acid concentration decreases.

Hence, pH = pKa + log {[salt] + 0.01 / [Acid] - 0.01}

pH = 4.76 + log {0.5 + 0.01/ 0.5 - 0.01}

pH = 4.76 + log (0.51/0.49)

pH = 4.76 + log (1.04)

= 4.76 + 0.017

pH = 4.77